The number of terms of the series 54,51,48,______so that their sum is 513?

Grade:11

1 Answers

Adithi
6 years ago
this is a simple AP where
a=54
d= – 3
sn= 513
 
by applying sn = n/2 [2a+(n-1)d] formula,
513 = n/2 [2*54 + (n-1)(– 3)]
n2 – 37n + 342 = 0
n =18 or 19
 
this implies that sum of 18 terms of the series = sum of 19 terms of the series = 513.
hope this helps. :)